A Bibliography of Publications about the AXIOM (formerly, Scratchpad) Symbolic Algebra Language

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 December 2023 Version 1.23

Title word cross-reference

1 [GJ71]. 1972 [Onl72]. 1984 [Fit84, GH84]. 1987 [Wex87].

5th [HP89].

6th [Mor89].

7th [SvADS07].

’84 [Fit84]. ’85 [BC85]. ’87 [Dav89, Wex87]. ’88 [Gia89]. ’89 [ACM89].

’90 [CC91, Mio90, WN90]. ’92 [Fit93, Wan92]. ’93 [Mio93]. ’95 [Lev95].

AAECC [HP89, Mor89]. AAECC-5 [HP89]. AAECC-6 [Mor89]. Abstract [JSW88, GL93, JSW86, Sch88, SW88]. abstraction [Boe89]. ACM [ACM89]. ACM-SIGSAM [ACM89]. Advances [Fat90]. advantages [MR90]. Algebra [BC85, Cal94, CGG^{+}04, DD05a, DST88, Dav89, DT90, DT92, DGT92, HP89, Jan88, Jen84a, Mor89, Sut85, Ben98, DGT91, DJ92, For90, GKW03, Joy08, Lam94, Mat89, Pag07, Roe95, SS88, Sme92, Web93]. Algebraic [ACM89, ACM94, Bro93, Bro07, DST88, DD89, Doo99, Fit84, GM89, Gia89, HP89, Lev95, Mor89, Pet71, Wan92, WN90, Wat91, Bro91, Duv95, Fat90, Rio92, vH94]. algebras [Kos91]. algébrique [Duv95]. algorithm [GM86, GM88, Sch89, Sit89, Sit92, vH94]. algorithme [Fou90]. Algorithmic [SDJ07]. Algorithms [DST88, HP89, Mor89]. American [Ano91]. Angeles [Pet71]. Annual [Ano91, Ano95]. Anwendung [Ben98]. Application [And88]. Applications [CC91, Lam94, Pur86, Ben98, GKW03]. applied [Dal92, Mat89, HP89, Mor89]. April [BC85, Mio90]. Argument [KN94]. Arithmetic [LM06]. asymptotic [Sal89, Sal91]. August [Bro07]. Austria [BC85, Mio93]. Automated [Doy99]. automatic [Sal89, Sal91]. automatics [Oll89]. autonomous [Sit89]. AXIOM [Ben98, GHK91, BT94, Dav92a, Dav92b, GS92, JS92, JT94, KN94, Ken01, Lam94, Sei94a, Sei94b, Zen92, Bou95, CGG^{+}04, Dal02, DD03, DD05c, DD05d, DD05e, DD05f, DD05h, DD05i, DD05a, DD05b, DD05g, Dal06, Doy99, Duv95, Gil92, JSD03, Joy08, Pag07, Pet93, PRDD07, Rio92, SDJ07].

Bad [Jan88]. balance [GBL91]. Bases [Arn95, GM86, GM89, Kos91, Sch91, SS88]. Basic [DT90, DT92]. basis [vH94]. BC [Doo99]. Beijing [SvADS07]. binary [ACS91]. birational [SS88]. Bonn [Wat91]. Browser [PRDD07]. Brunel [Onl72]. Buchberger [GM86, GM88]. Bundespost [GHK91].

calcul [SSV87]. Calculations [BT94]. calculus [Ken01]. California [Pet71]. Cambridge [Fit84]. Canada [Doo99, Bro07, Lev95]. canonical [Gil92]. Capri [Mio90]. capsule [GJ72b]. Carmichael [Arn95]. Categorical [DGT92, DGT91]. Challenges [Ano91]. Changing [Ano91]. ChE [LeB91]. China [SvADS07]. clôture [Duv95]. class [Wan91]. classroom [LeB91]. closure [DD89, Duv95, Rio92]. Codes [HP89, Mor89, ACS91]. Coding [CC91]. Coercion [SJ87a, Doy99, For90, SJ87b]. coherence [Web93]. combinatoire [CCM92]. combinatorial [CCM92]. combinatorics [And88]. Command [DD05f]. Commutative [DT90, DT92]. Compiler [DD05i, Jen76]. Completion [Sei94a]. Computation [ACM89, Bee17, Bro93, Bro07, DST88, Fit84, Fit93, Gia89, Gil92, JOS93, JSW88, JS92, JSD03, Lev95, LMS07, Mio90, Mio93, Wan92, Wat91, Dal92, JSW86, Mon93, SW88, Wat89, ACM94, Doo99, WN90]. computational [Mat89, SvADS07]. Computer [BC85, Cal94, CGG^{+}04, DST88, Dav89, GKW03, Jan88, Jen84a, Sut85, Ben98, DJ92, For90, Joy08, Lam94, LM91, MR90, Pag07, Roe95, Sme92, Web93]. Computeralgebra [Ben98, GHK91]. Computeralgebra-System [GHK91]. Computeralgebra-Systemen [Ben98]. Computing [Nor75, Onl72, Wan89, vH94]. concept [Lam94]. concepts [GBL91]. Conference [Ano91, Ano92, BC85, Dav89, GH84, HP89, Mor89, SvADS07, Onl72]. constants [Wan89]. Constructing [Arn95]. construction [Fat90]. Correcting [HP89, Mor89]. Crystal [DD05b]. cubical [BT94]. curve [Bro91].

dans [CCM92]. data [GL93, Sch88]. Datatype [JSW88, SW88, JSW86]. Decomposition [BS93]. DERIVE [Ben98]. design [Fat90, Fit93, Mio90, Mio93]. determine [SS88]. Deutschen [GHK91]. Developer [CGG^{+}04]. Developers [DD05e]. development [GL93]. Developments [JOS93]. differential [Bro91, Sch89, Sei94b, Wan91]. Differentialformen [Zen92]. differentiation [SDJ07]. Dimick [CGG^{+}04]. DISCO [Fit93, Mio90, Mio93]. distance [ACS91]. Does [Dav92b]. domain [Mon93]. Double [BS94]. Double-track [BS94]. d’un [SSV87]. Dynamic [BGDW95, Duv95]. dynamique [Duv95].

Education [Ano91]. Efficient [For90, LM06]. electronic [LM91]. elementary [Bro89]. energy [GBL91]. Engineering [Ano91, Ben98]. England [Fit84, Onl72]. English [CCM92]. Environment [LM06]. Environments [Ano92, Pur86]. equation [Bro91]. equations [GM89, Sch89]. equivalence [SS88]. Error [HP89, Mor89]. Error-Correcting [HP89, Mor89]. EUROCAL [BC85, Dav89]. EUROCODE [CC91]. European [BC85, Dav89]. EUROSAM [Fit84]. Evaluation [BGDW95, Duv95, Duv95]. exact [Roe95]. Examples [DJ92, Sal89, Sal91, KN94]. expansions [Sal89, Sal91]. Experience [GJ72a, Pet93]. Extension [Jen71]. extensions [DD89].

Facilities [SJ87a, SJ87b]. Facility [Jen71, GJ71]. Factorization [DGT92, DGT91, Luc86, Sch89]. fast [Luc86]. field [Rio92, vH94]. Fields [GS92]. Finite [GS92]. first [Sit89]. first-order [Sit89]. fixed [Wat89]. form [Gil92]. Formal [Nor75]. formel [SSV87]. Formelmanipulationssystemes [Lue77]. formulas [Kos91]. Forschungsinstituts [GHK91]. foundations [GKW03]. Fraction [BS93]. Fraser [Doo99]. free [Kos91]. French [Duv95]. Full [BS93]. Function [Bee17, vH94]. functional [Dal92]. Functions [BS93, And88, Bro89, Mon93, Wan89]. future [BS94].

gain [BS94]. games [Kos91]. GAMM [Ano95]. Gauss [GM94, Mon93]. generation [Lam94]. Generic [LMS07]. geometry [KKM89]. German [Ben98]. Germany [Wat91]. Gmunden [Mio93]. Goldman [Sit89]. graded [GL93]. graphes [CCM92]. Graphics [DD05h]. graphs [CCM92]. Gröbner [GM89, Sch91, SS88].

Gröbnerbasen [Zen92]. groups [BT94]. Guide [DD05c, DD05d, DD05e].

Hamming [CCM92]. Handbook [Bee17, GKW03]. High [Ano92]. High-Level [Ano92]. homological [Lam91]. homomorphisms [SS88]. Hyperdoc [DD05g].

ICCS [SvADS07]. identifiability [Oll89]. identities [ACS91, Bur91]. IFIP [Ano92]. ihre [Zen92]. II [BW87, BW89, DGT91, DGT92, DD89, JSW86, JSW88, KKM89, SSV87, SvADS07, Sut85, SJ87b, SJ87a, SW88, Wan89]. IMACS [JOS93]. Implantation [Fou90]. Implementation [BGDW95, DD05a, Fit93, KKM89, LM06, Mio90, Mio93, Luc86, Rio92]. Implementierung [Lue77, Zen92]. inexpensive [LM91]. Inference [SJ87a, Boe89, For90, SJ87b]. Infinite [BW87, BW89]. Ingenieurmathematik [Ben98]. installation [GM88]. integrable [Sei94b]. integral [vH94]. interactive [GJ71, Onl72]. interconnection [Pur86]. International [ACM89, ACM94, Bro93, Bro07, CC91, Doo99, Fit84, Fit93, Gia89, HP89, JOS93, Jan88, Lev95, Mio90, Mio93, Mor89, SvADS07, WN90, Wat91, Onl72, Wan92]. Interpreter [DD03, SJ87a, SJ87b]. Interpreters [Wex87]. Interpretive [Wex87]. Introduction [GM94]. Inversibility [Oll89]. involution [Sei94a]. involving [KN94]. ISSAC [ACM89, Gia89, Doo99, Lev95, WN90]. ISSAC’91 [Wat91]. ISSAC’93 [Bro93]. ISSAC’94. [ACM94]. Itô [Ken01]. Italy [Gia89, Mio90, Mor89]. iteration [Mat89].

Jenks [CGG^{+}04]. Jordan [Gil92]. July [Bro93, Bro07, Doo99, Fit84, Gia89, GH84, Lev95, Mor89, Wat91]. June [Wex87]. keys [Jen84b]. Kiev [Bro93]. Kovacic [Fou90]. kryptologischen [GHK91].

langage [SSV87]. Language [Jen84a, Sut85, Dal92, Gil92, Jen74]. Level [Ano92, LM06]. Liapunov [Wan89]. Library [Bee17]. Lie [Kos91]. linear [GM86, Sch89, Sit92]. link [KN94]. Linz [BC85]. liquids [Roe95].

MACSYMA [Ben98, GH84]. make [JT94]. Manipulation [Pet71, Fat90, Wan91]. MAPLE [Ben98]. mappings [Oll89]. March [Pet71]. material [GBL91]. math [GJ72a]. MATHCAD [Ben98, BS94, GBL91, LeB91, MR90]. MathCW [Bee17]. MATHEMATICA [Ben98]. Mathematical [Bee17, JSW88, JSW86, Pur86, SW88]. Mathematical-Function [Bee17]. mathematics [Ben98, GJ71]. MATLAB [Ben98]. matrix [Gil92]. May [Jan88, SvADS07]. Mechanical [Wan91, LM91]. Meeting [Ano95]. membership [SS88]. Memoriam [CGG^{+}04]. META [Jen71]. META/PLUS [Jen71]. method [Bou95, Wat89]. minimum [ACS91]. Minnesota [Wex87]. Monomial [Sch91]. Montréal [Lev95]. multinomial [Sit89]. Multiple [LM06]. Multiple-Level [LM06]. MuPAD [Ben98]. MuPAD. [Ben98].

NAG [KN94]. Neuenahr [Jan88]. Newton [ACS91]. Newtonian [Roe95]. Next [Lam94]. non [Roe95]. non-Newtonian [Roe95]. nonlinear [KN94]. nouveau [SSV87]. Numbers [Arn95].

Object [Bou95, Sme92]. One [Dav92b]. ongoing [Ken01]. online [GJ72a, Onl72, Onl72]. Ontario [Bro07]. Open [Dal02, Joy08, Pag07]. operators [Sei94b]. order [Sit89]. ordered [Rio92]. orderings [Sch91]. ordinary [Sch89]. oriented [Bou95, Sme92].

package [KKM89, Sch88]. parameterized [Mon93]. parameters [Roe95]. parametric [Sit92]. Part [SvADS07]. Partial [BS93]. pattern [Jen76]. Paul [Wex87]. perturbation [Lam91, LM91]. Picard [Mat89]. pictorial [Sme92]. Pioneer [CGG^{+}04]. PLUS [Jen71, BS94]. point [Wat89]. polymorphic [Dal92]. Polynomial [LM06, GM89, Luc86]. Portable [Bee17]. Power [Nor75, Wat89]. presence [Boe89]. Présentation [SSV87]. Primality [Dav93]. primer [Jen84b]. Problem [Ano92, CCM92, DJ92, Pur86]. problème [CCM92]. problems [And88]. Proceedings [ACM94, Lev95, SvADS07, WN90, Bro93, Doo99, Onl72, ACM89, Ano91, Bro07, Cal94, CC91, Dav89, Fit93, Gia89, GH84, HP89, JOS93, Jan88, Mio90, Mio93, Mor89, Pet71, Wat91, Wex87, BC85]. Production [KN94]. Program [Dav92b, Wan89]. Programmers [DD05d]. Programming [Ano92, Bee17, LM06, LMS07, Sch88, Gil92, Sme92]. proving [KKM89]. Pseudo [Sei94b]. Pseudoprimes [Arn95].

quantum [LM91].

Ramanujan [And84, Bur91]. Rational [BS93, Oll89]. Real [Rio92, Bro89]. research [Lam94]. Resolutions [Lam91]. Revisited [Dav93]. Rhine [Cal94]. Richard [CGG^{+}04]. rings [GL93]. Risch [Bro91]. Rogers [Bur91]. Rome [Gia89, Mor89].

sa [CCM92]. Schenectady [GH84]. Science [SvADS07]. Scientific [Ano92, JS92, JSD03]. Scratchpad [JSW86, SSV87, LM91, And84, BW87, GJ71, GJ72a, GJ72b, Jen71, Jen74, Jen84a, KKM89, And88, BW89, Bur91, CCM92, DGJ^{+}84, DT90, DGT91, DT92, DGT92, DD89, Fou90, Jen84b, JSW88, JT94, Lam94, Sch88, Sut85, SJ87b, SJ87a, SW88, Wan89]. SCRATCHPAD/1 [GJ71]. second [Pet71]. September [Mio93, Onl72]. Series [Nor75, Wat89]. Several [Arn95, Kos91]. signature [Mon93]. SIGPLAN [Wex87]. SIGSAM [ACM89]. Simon [Doo99]. simplicial [BT94]. Simplification [Bro89]. Society [Ano91]. Software [Bee17, MR90, Pur86]. solution [CCM92, GM89, Roe95]. solutions [Roe95]. Solving [Ano92, DJ92, Pur86, Sit89, Sit92]. some [ACS91]. Source [Dal02, Joy08, Pag07]. SPDE [Sch88]. special [And88]. split [SS88]. square [Gil92]. St [Wex87]. staggered [GM86]. Standard [BS94]. story [Ken01]. Strong [Arn95]. structural [Oll89]. Structures [BW87, BW89]. SubPrograms [KN94]. support [Mon93]. surjective [SS88]. Symbolic [ACM89, ACM94, Bro93, Bro07, Doo99, Fit84, Fit93, JOS93, Ken01, Lev95, LMS07, Mat89, Mio90, Mio93, Pet71, Wan92, WN90, Wat91, Dal92, GJ71, GJ72a, Gia89]. symmetry [Sch88]. Symposium [ACM89, ACM94, Bro93, Bro07, CC91, Doo99, Fit84, Fit93, Gia89, JOS93, Jan88, Lev95, Mio90, Mio93, Wan92, WN90, Wat91, Wex87, Pet71]. Syntax [Jen71]. System [Dav92a, Dav92b, GHK91, Jen84a, JSW88, JS92, JSD03, Sut85, Wan92, GJ72a, JSW86, Mon93, Pag07, Pur86, SW88]. Systemen [Ben98]. Systems [DST88, Fit93, Mio90, Mio93, Ben98, Fat90, GM89, GKW03, Joy08, KN94, Lam94, Sei94b, Sit89, Sit92, Wan91].

TC2 [Ano92]. TC2/WG [Ano92]. Teaching [MR90, GBL91]. technique [LM91]. Techniques [Wex87]. TELEKOM [GHK91]. Testing [Dav93]. theorem [KKM89]. Theorist [LeB91]. Theory [CC91]. track [BS94]. transparent [MR90].

Trends [JOS93, Fat90, Jan88]. Tutorial [Dal06]. Type [Boe89, SJ87a, For90, GL93, SJ87b]. types [Sch88].

Ueberlegungen [Lue77]. Ukraine [Bro93]. University [Bro07, Doo99, Onl72]. Untersuchungen [GHK91]. use [LeB91]. User [GH84]. Users [DD05c, BS94]. Using [Bee17, GBL91, SS88, Sme92, DJ92, GM89, Gil92, LM91, Roe95]. Uxbridge [Onl72].

Vancouver [Doo99]. Verified [Roe95].

verkehrstheoretischen [GHK91]. versions [BS94]. via [ACS91, Lam91]. View [DT90, DT92, DGT92, DGT91, GJ72b]. Virtues [LMS07]. Volume [DD03, DD05c, DD05d, DD05e, DD05f, DD05h, DD05i, DD05a, DD05b, DD05g, Dal06, JSD03, PRDD07].

Waterloo [Bro07]. WG [Ano92]. Which [Arn95]. will [BS94]. Word [Kos91]. Working [Ano92]. Workshop [Cal94]. World [Ano91].

York [GH84].

zur [Lue77].

References

ACM:1989:PAI

[ACM89] ACM, editor. Proceedings of the ACM-SIGSAM 1989 International Symposium on Symbolic and Algebraic Computation, ISSAC ’89. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

ACM:1994:IPI

[ACM94] ACM, editor. ISSAC’94. Proceedings of the International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994.

Augot:1991:MDS

[ACS91] D. Augot, P. Charpin, and N. Sendrier. The minimum distance of some binary codes via the Newton’s identities. In Cohen and Charpin [CC91], pages 65–73. ISBN 0-387-54303-1 (New York), 3-540-54303-1 (Berlin). LCCN QA268.E95 1990.

Andrews:1984:RS

[And84] George E. Andrews. Ramanujan and SCRATCHPAD. In Golden and Hussain [GH84], pages 383–??

Andrews:1988:ASP

[And88] G. E. Andrews. Application of Scratchpad to problems in special functions and combinatorics. In Janßen [Jan88], pages 158–?? ISBN 3-540-18928-9, 0-387-18928-9. LCCN QA155.7.E4T74 1988.

Anonymous:1991:PAC

[Ano91] Anonymous, editor. Proceedings 1991 Annual Conference, American Society for Engineering Education. Challenges of a Changing World. ASEE, Washington, DC, USA, 1991. 2 vol.

Anonymous:1992:PEH

[Ano92] Anonymous. Programming environments for high-level scientific problem solving. IFIP TC2/WG 2.5 working conference. IFIP Transactions. A. Computer Science and Technology, A-2:??, 1992. CODEN ITATEC. ISSN 0926-5473.

Anonymous:1995:GAM

[Ano95] Anonymous. GAMM 94 annual meeting. Zeitschrift fur Angewandte Mathematik und Physik, 75(suppl. 2), 1995. CODEN ZAMMAX. ISSN 0044-2267.

Arnault:1995:CCN

[Arn95] François Arnault. Constructing Carmichael numbers which are strong pseudoprimes to several bases. Journal of Symbolic Computation, 20(2):151–162 (or 151– 161??), August 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Buchberger:1985:EEC

[BC85] Bruno Buchberger and Bob F. Caviness, editors. EUROCAL ’85: European Conference on Computer Algebra, Linz, Austria, April 1–3, 1985: proceedings, volume 204 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-15983-5 (vol. 1), 0-387-15984-3 (vol. 2). LCCN QA155.7.E4 E86 1985. Two volumes.

Beebe:2017:MFC

[Bee17] Nelson H. F. Beebe. The Mathematical-Function Computation Handbook: Programming Using the MathCW Portable Software Library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-64109-3 (hardcover), 3-319-64110-7 (e-book). xxxvi + 1114 pp. LCCN QA75.5-76.95.

URL http://www.springer.com/
us/book/9783319641096.

Benker:1998:ICS

[Ben98] Hans Benker. Ingenieurmathematik mit Computeralgebra-Systemen. AXIOM, DERIVE, MACSYMA, MAPLE, MATHCAD, MATHE-MATICA, MATLAB und MuPAD in der Anwendung. (German) Engineering mathematics with computer algebra systems. The applications: AXIOM, DERIVE, MACSYMA, MAPLE, MATHCAD, MATHE-MATICA, MATLAB UND Mu-PAD. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1998. xiii + 439 pp.

Broadbery:1995:IDE

[BGDW95] P. A. Broadbery, T. Gómez-Díaz, and S. M. Watt. On the implementation of dynamic evaluation. In Levelt [Lev95], pages 77–84. ISBN 0-89791-699-

9. LCCN QA 76.95 I59 1995.

URL http://www.acm.org:80/
pubs/citations/proceedings/issac/
220346/p77-broadbery/. ACM or-

der number: 505950.

Boehm:1989:TIP

[Boe89] Hans-J. Boehm. Type inference in the presence of type abstraction. ACM SIGPLAN Notices, 24 (7):192–206, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.acm.
org:80/pubs/citations/proceedings/
pldi/73141/p192-boehm/.

Boulanger:1995:OOM

[Bou95] J.-L. Boulanger. Object oriented method for Axiom. ACM SIG-PLAN Notices, 30(2):33–41, February 1995. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bronstein:1989:SRE

[Bro89] M. Bronstein. Simplification of real elementary functions. In ACM [ACM89], pages 207–211. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

Bronstein:1991:RDE

[Bro91] M. Bronstein. The Risch differential equation on an algebraic curve. In Watt [Wat91], pages 241–246. ISBN 0-89791-437-6. LCCN QA76.95.I59 1991.

Bronstein:1993:IPI

[Bro93] Manuel Bronstein, editor. IS-SAC’93: proceedings of the 1993 International Symposium on Symbolic and Algebraic Computation, July 6– 8, 1993, Kiev, Ukraine. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993. ACM order number: 505930.

Brown:2007:PIS

[Bro07] C. W. Brown, editor. Proceedings of the 2007 International Symposium on Symbolic and Algebraic Computation, July 29–August 1, 2007, University of Waterloo, Waterloo, Ontario, Canada. ACM Press, New York, NY 10036, USA, 2007. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Bronstein:1993:FPF

[BS93] Manuel Bronstein and Bruno Salvy. Full partial fraction decomposition of rational functions. In Bronstein [Bro93], pages 157–160. ISBN 0-89791-604-2. LCCN QA 76.95 I59 1993.

URL http://www.acm.org:80/
pubs/citations/proceedings/issac/
164081/p157-bronstein/. ACM

order number: 505930.

Beneke:1994:DFM

[BS94] T. Beneke and W. Schwippert.

Double-track into the future: Math-CAD will gain new users with Standard and Plus versions. Elektronik, 43(15):107–110, July 1994. CODEN EKRKAR. ISSN 0013-5658.

Brown:1994:CSC

[BT94] R. Brown and A. Tonks. Calculations with simplicial and cubical groups in AXIOM. Journal of Symbolic Computation, 17(2):159–179, February 1994. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Burge:1991:SRI

[Bur91] W. H. Burge. Scratchpad and the Rogers–Ramanujan identities. In Watt [Wat91], pages 189–190. ISBN 0-89791-437-6. LCCN QA76.95.I59 1991.

Burge:1987:ISS

[BW87] W. Burge and S. Watt. Infinite structures in SCRATCHPAD II. Technical Report RC 12794 (#57573), IBM Thomas J. Watson Research Center, Bos 218, Yorktown Heights, NY 10598, USA, 1987. ?? pp.

Burge:1989:ISS

[BW89] W. H. Burge and S. M. Watt. Infinite structures in Scratchpad II. In Davenport [Dav89], pages 138– 148. ISBN 3-540-51517-8. LCCN QA155.7.E4E86 1987.

Calmet:1994:RWC

[Cal94] J. Calmet, editor. Rhine Workshop on Computer Algebra. Proceedings. Universität Karlsruhe, Karlsruhe, Germany, 1994.

Cohen:1991:EIS

[CC91] G. Cohen and P. Charpin, editors. EUROCODE ’90. International Symposium on Coding Theory and Applications Proceedings. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-54303-1 (New York), 3-540-54303-1 (Berlin). LCCN QA268.E95 1990.

Camion:1992:PCG

[CCM92] Paul Camion, Bernard Courteau, and Andre Montpetit. Un problème combinatoire dans les graphes de Hamming et sa solution en Scratchpad. (English: A combinatorial problem in Hamming graphs and its solution in Scratchpad). Rapports de recherche 1586, Institut National de Recherche en Informatique et en Automatique, Le Chesnay, France, January 1992. 12 pp.

Caviness:2004:MRD

[CGG^{+}04] Bob Caviness, Barbara Gatje, James H. Griesmer, Tony Hearn, Manual Bronstein, and Erich Kaltofen. In memoriam: Richard Dimick Jenks: Axiom developer and computer algebra pioneer. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 38(1):30, March 2004. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (elec-

tronic). URL http://savannah.
nongnu.org/projects/axiom/.

Dalmas:1992:PFL

[Dal92] S. Dalmas. A polymorphic functional language applied to symbolic computation. In Wang [Wan92], pages 369–375. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992.

Daly:2002:AOS

[Dal02] T. Daly. Axiom as open source.

SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 36(1):28–??, March 2002. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Daly:2006:AVA

[Dal06] Timothy Daly. Axiom Volume 1:

Axiom Tutorial. Lulu, Inc., 860 Aviation Parkway, Suite 300, Morrisville, NC 27560, USA, 2006. ISBN 1-4116-6597-X. iv + 285 pp. LCCN ???? US$15.00. URL http://www.

axiom-developer.org/axiom-website/
bookvol1.pdf;   http://www.lulu.
com/content/190827.

Davenport:1989:EEC

[Dav89] J. H. Davenport, editor. EUROCAL ’87. European Conference on Computer Algebra Proceedings. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-51517-8. LCCN QA155.7.E4E86 1987.

Davenport:1992:AS

[Dav92a] J. H. Davenport. The AXIOM system. AXIOM Technical Report TR5/92 (ATR/3) (NP2492), Numerical Algorithms Group, Inc., Downer’s Grove, IL, USA and Oxford, UK, December 1992. ?? pp.

                                        URL http://www.nag.co.uk/doc/
                                        TechRep/axiomtr.html.

Davenport:1992:HDO

[Dav92b] J. H. Davenport. How does one program in the AXIOM system? AXIOM Technical Report TR6/92 (ATR/4) (NP2493), Numerical Algorithms Group, Inc., Downer’s Grove, IL, USA and Oxford, UK, December 1992. ?? pp.

                                        URL http://www.nag.co.uk/doc/
                                        TechRep/axiomtr.html.

Davenport:1993:PTR

[Dav93] J. H. Davenport. Primality testing revisited. Technical Report TR2/93 (ATR/6) (NP2556), Numerical Algorithms Group, Inc., Downer’s Grove, IL, USA and Oxford, UK, August 1993. ?? pp.

                                        URL http://www.nag.co.uk/doc/
                                        TechRep/axiomtr.html.

Dicrescenzo:1989:AEA

[DD89] C. Dicrescenzo and D. Duval. Algebraic extensions and algebraic closure in Scratchpad II. In Gianni [Gia89], pages 440–446. ISBN 3-540-51084-2. LCCN QA76.95 .I57 1988. Conference held jointly with AAECC-6.

Daly:2003:AVA

[DD03] Tim Daly and Martin Dunstan. Axiom Volume 5: Axiom Interpreter. 2003. xlvi + 1387 pp. LCCN

????  URL http://www.axiom-
developer.org/axiom-website/
bookvol5.pdf.

Daly:2005:AVAh

[DD05a] Tim Daly and Martin Dunstan. Axiom Volume 10: Axiom Algebra: Implementation.

2005.  iv + 5 pp.  LCCN

????  URL http://www.axiom-
developer.org/axiom-website/
bookvol10.pdf.

Daly:2005:AVAi

[DD05b] Tim Daly and Martin Dunstan.

Axiom Volume 12: Axiom Crystal. 2005. iv + 9 pp. LCCN

????  URL http://www.axiom-
developer.org/axiom-website/
bookvol12.pdf.

Daly:2005:AVAb

[DD05c] Tim Daly and Martin Dunstan.

Axiom Volume 2: Axiom Users Guide. 2005. iv + 7 pp. LCCN

????  URL http://www.axiom-
developer.org/axiom-website/
bookvol2.pdf.

Daly:2005:AVAc

[DD05d] Tim Daly and Martin Dunstan. Axiom Volume 3: Axiom Programmers Guide. 2005. iv + 3 pp. LCCN

                                        ????  URL http://www.axiom-
                                        developer.org/axiom-website/
                                        bookvol3.pdf.

Daly:2005:AVAd

[DD05e] Tim Daly and Martin Dunstan. Axiom Volume 4: Axiom Developers Guide. 2005. v + 91 pp. LCCN

                                        ????  URL http://www.axiom-
                                        developer.org/axiom-website/
                                        bookvol4.pdf.

Daly:2005:AVAe

[DD05f] Tim Daly and Martin Dunstan. Axiom Volume 6: Axiom Command. 2005. vi + 187 pp. LCCN

                                        ????  URL http://www.axiom-
                                        developer.org/axiom-website/
                                        bookvol6.pdf.

Daly:2005:AVAj

[DD05g] Tim Daly and Martin Dunstan. Axiom Volume 7: Axiom Hyperdoc. 2005. xvi + 632 pp. LCCN

                                        ????  URL http://www.axiom-
                                        developer.org/axiom-website/
                                        bookvol7.pdf.

Daly:2005:AVAf

[DD05h] Tim Daly and Martin Dunstan. Axiom Volume 8: Axiom Graphics. 2005. xi + 538 pp. LCCN

                                        ????  URL http://www.axiom-
                                        developer.org/axiom-website/
                                        bookvol8.pdf.

Daly:2005:AVAg

[DD05i] Tim Daly and Martin Dunstan. Axiom Volume 9: Axiom Compiler. 2005. iv + 30 pp. LCCN

                                        ????  URL http://www.axiom-
                                        developer.org/axiom-website/
                                        bookvol9.pdf.

Davenport:1984:S

[DGJ^{+}84] J. Davenport, P. Gianni, R. Jenks, V. Miller, S. Morrison, M. Rothstein, C. Sundaresan, R. Sutor, and B. Trager. Scratchpad. Mathematical Sciences Department, IBM Thomas Watson Research Center, 1984. ?? pp.

Davenport:1991:SVA

[DGT91] J. H. Davenport, P. Gianni, and B. M. Trager. Scratchpad’s view of algebra. II. A categorical view of factorization. In Watt [Wat91], pages 32–38. ISBN 0-89791-437-6. LCCN QA76.95.I59 1991. auch in: AXIOM Technical Report, ATR/2, NAG Ltd., Oxford, 1992.

Davenport:1992:SVAb

[DGT92] J. H. Davenport, P. Gianni, and B. M. Trager. Scratchpad’s view of algebra II: A categorical view of factorization. Technical Report TR4/92 (ATR/2) (NP2491), Numerical Algorithms Group, Inc., Downer’s Grove, IL, USA and Oxford, UK, December 1992. ?? pp.

URL http://www.nag.co.uk/doc/
TechRep/axiomtr.html.

Duval:1992:EPS

[DJ92] D. Duval and F. Jung. Examples of problem solving using computer algebra. IFIP Transactions. A. Computer Science and Technology, A-2:133–141, 143, 1992. CO-DEN ITATEC. ISSN 0926-5473.

Dooley:1999:IJS

[Doo99] Sam Dooley, editor. ISSAC 99: July 29–31, 1999, Simon Fraser University, Vancouver, BC, Canada: proceedings of the 1999 International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 1999. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999.

Doye:1999:ACA

[Doy99] Nicolas J. Doye. Automated coercion for Axiom. In Dooley [Doo99], pages 229–235. ISBN 1-58113-073-2. LCCN QA76.95 .I57 1999.

URL http://www.acm.org/pubs/
contents/proceedings/issac/309831/

.

Davenport:1988:CA

[DST88] J. H. Davenport, Y. Siret, and E. Tournier. Computer Algebra: Systems and Algorithms for Algebraic Computation. Academic Press, New York, NY, USA, 1988. ISBN 0-12-204230-1. xix + 267 pp. LCCN QA155.7.E4 D38 1988.

Davenport:1990:SVA

[DT90] J. H. Davenport and B. M. Trager. Scratchpad’s view of algebra I: Basic commutative algebra. In Miola [Mio90], pages 40–54. ISBN 0-387-52531-9 (New York), 3-540-52531-9 (Berlin). LCCN QA76.9.S88I576 1990. auch in: AXIOM Technical Report, ATR/1, NAG Ltd., Oxford, 1992.

Davenport:1992:SVAa

[DT92] J. H. Davenport and B. M. Trager. Scratchpad’s view of algebra I: Basic commutative algebra. Technical Report TR3/92 (ATR/1) (NP2490), Numerical Algorithms Group, Inc., Downer’s Grove, IL, USA and Oxford, UK, December 1992. ?? pp.

URL http://www.nag.co.uk/doc/
TechRep/axiomtr.html.

Duval:1995:DEA

[Duv95] Dominique Duval. Évaluation dynamique et clôture algébrique en Axiom. (french) Dynamic evaluation and algebraic closure in Axiom. J. Pure Appl. Algebra, 99(3):267– 295, 1995. ISSN 0022-4049.

Fateman:1990:ATD

[Fat90] R. J. Fateman. Advances and trends in the design and construction of algebraic manipulation systems. In Watanabe and Nagata [WN90], pages 60–67. ISBN 0-89791-401-5. LCCN QA76.95 .I57 1990.

Fitch:1984:E

[Fit84] J. P. Fitch, editor. EUROSAM ’84: International Symposium on Symbolic and Algebraic Computation, Cambridge, England, July 9– 11, 1984, volume 174 of Lecture Notes In Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 0-387-13350-X. LCCN QA155.7.E4 I57 1984.

Fitch:1993:DIS

[Fit93] J. Fitch, editor. Design and Implementation of Symbolic Computation Systems International Symposium, DISCO ’92 Proceedings. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-57272-4 (New York), 3-540-57272-4 (Berlin). LCCN QA76.9.S88I576 1992.

Fortenbacher:1990:ETI

[For90] A. Fortenbacher. Efficient type inference and coercion in computer algebra. In Miola [Mio90], pages 56–60. ISBN 0-387-52531-9 (New York), 3-540-52531-9 (Berlin). LCCN QA76.9.S88I576 1990.

Fouche:1990:ILK

[Fou90] Francois Fouche. Une implantation de l’algorithme de Kovacic en Scratchpad. Technical report, Institut de Recherche Mathématique Avancée, Strasbourg, France, 1990. 31 pp.

Goodwin:1991:UMT

[GBL91] B. M. Goodwin, R. A. Buonopane, and A. Lee. Using MathCAD in teaching material and energy balance concepts. In Anonymous [Ano91], pages 345–349 (vol. 1). 2 vol.

Golden:1984:PMU

[GH84] V. Ellen Golden and M. A. Hussain, editors. Proceedings of the 1984 MACSYMA Users’ Conference: Schenectady, New York, July 23–25, 1984. General Electric, Schenectady, NY, USA, 1984.

Grabmeier:1991:CSA

[GHK91] J. Grabmeier, K. Huber, and U. Krieger. Das Computeralgebra-System AXIOM bei kryptologischen und verkehrstheoretischen Untersuchungen des Forschungsinstituts der Deutschen Bundespost TELEKOM. Technischer Report TR 75.91.20, IBM Wissenschaftliches Zentrum, Heidelberg, Germany, 1991. ?? pp.

Gianni:1989:SAC

[Gia89] P. (Patrizia) Gianni, editor. Symbolic and Algebraic Computation. International Symposium ISSAC ’88, Rome, Italy, July 4–8, 1988. Proceedings, volume 358 of Lecture Notes In Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-51084-2. LCCN QA76.95 .I57 1988. Conference held jointly with AAECC-6.

Gil:1992:CJC

[Gil92] I. Gil. Computation of the Jordan canonical form of a square matrix (using the Axiom programming language). In Wang [Wan92], pages 138–145. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992.

Griesmer:1971:SIF

[GJ71] J. H. Griesmer and R. D. Jenks.

SCRATCHPAD/1 — an interactive facility for symbolic mathematics. In Petrick [Pet71], pages 42–58. LCCN QA76.5 .S94 1971.

URL http://delivery.acm.org/
10.1145/810000/806266/p42-griesmer.
pdf.

Griesmer:1972:EOSb

[GJ72a] J. Griesmer and R. Jenks. Experience with an online symbolic math. system SCRATCHPAD. In Online’72 [Onl72], page ?? ISBN 0-903796-02-3. LCCN QA76.55 .O54 1972. Two volumes.

Griesmer:1972:SCV

[GJ72b] James H. Griesmer and Richard D.

Jenks. SCRATCHPAD: A capsule view. ACM SIGPLAN Notices, 7(10):93–102, 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Proceedings of the symposium on Two-dimensional manmachine communication, Mark B. Wells and James B. Morris (eds.).

Grabmeier:2003:CAH

[GKW03] Johannes Grabmeier, Erich Kaltofen, and Volker Weispfenning, editors. Computer algebra handbook: foundations, applications, systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-65466-6. xx + 637 pp. LCCN QA155.7.E4 C64954 2003.

URL http://www.springer.com/
sgw/cda/frontpage/0,11855,1-
102-22-1477871-0,00.html. In-

cludes CD-ROM.

Goodloe:1993:ADT

[GL93] A. Goodloe and P. Loustaunau. An abstract data type development of graded rings. In Fitch [Fit93], pages 193–202. ISBN 0-387-57272-4 (New York), 3-540-57272-4 (Berlin). LCCN QA76.9.S88I576 1992.

Gebauer:1986:BAS

[GM86] Rüdiger Gebauer and H. Michael Möller. Buchberger’s algorithm and staggered linear bases. In Bruce W. Char, editor, Proceedings of the 1986 Symposium on Symbolic and Algebraic Computation: Symsac ’86, July 21–23, 1986, Waterloo, Ontario, pages 218–221. ACM Press, New York, NY 10036, USA, 1986. ISBN 0-89791-199-7. LCCN QA155.7.E4 A281 1986. ACM order number 505860.

Gebauer:1988:IBA

[GM88] R. Gebauer and H. M. Möller. On an installation of Buchberger’s algorithm. Journal of Symbolic Computation, 6(2-3):275–286, October– December 1988. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Gianni:1989:ASS

[GM89] P. Gianni and T. Mora. Algebraic solution of systems of polynomial equations using Gröbner bases. In Huguet and Poli [HP89], pages 247– 257. ISBN 3-540-51082-6. LCCN QA268.A35 1987.

Gruntz:1994:IG

[GM94] D. Gruntz and M. Monagan. Introduction to Gauss. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 28(2):3–19, August 1994. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Grabmeier:1992:FFA

[GS92] J. Grabmeier and A. Scheerhorn.

Finite fields in AXIOM. AXIOM Technical Report TR7/92 (ATR/5) (NP2522), Numerical Algorithms Group, Inc., Downer’s Grove, IL, USA and Oxford, UK, December 1992. ?? pp. URL http:

//www.nag.co.uk/doc/TechRep/
axiomtr.html.

Huguet:1989:AAA

[HP89] L. Huguet and A. Poli, editors. Applied Algebra, Algebraic Algorithms and Error-Correcting Codes. 5th International Conference, AAECC-5 Proceedings. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-51082-6. LCCN QA268.A35 1987.

Janssen:1988:TCA

[Jan88] R. Janßen, editor. Trends in Computer Algebra, International Symposium Bad Neuenahr, May 19– 21, 1987, Proceedings, volume 296 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 3-540-18928-9, 0-387-18928-9. LCCN QA155.7.E4T74 1988.

Jenks:1971:MPS

[Jen71] R. D. Jenks. META/PLUS: The syntax extension facility for SCRATCHPAD. Research Report RC 3259, International Business Machines Inc., Thomas J. Watson Research Center, Yorktown Heights, NY, USA, February 1971. ?? pp.

Jenks:1974:SL

[Jen74] R. D. Jenks. The SCRATCHPAD language. ACM SIGPLAN Notices, 9(4):101–111, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jenks:1976:PC

[Jen76] Richard D. Jenks. A pattern compiler. In Richard D. Jenks, editor, Symsac ’76: proceedings of the 1976 ACM Symposium on Symbolic and Algebraic Computation, August 10– 12, 1976, Yorktown Heights, New York, pages 60–65. ACM Press, New York, NY 10036, USA, 1976. ISBN ???? LCCN QA155.7.E4 .A15 1976; QA9.58 .A11 1976.

Jenks:1984:NSL

[Jen84a] Richard D. Jenks. The new SCRATCHPAD language and system for computer algebra. In Golden and Hussain [GH84], pages 409–??

Jenks:1984:PKN

[Jen84b] Richard D. Jenks. A primer: 11 keys to New Scratchpad. In Fitch [Fit84], pages 123–147. ISBN 0-387-13350-X. LCCN QA155.7.E4 I57 1984.

Jacob:1993:PSI

[JOS93] G. Jacob, N. E. Oussous, and S. Steinberg, editors. Proceedings SC 93. International IMACS Symposium on Symbolic Computation. New Trends and Developments. LIFL Univ. Lille, Lille, France, 1993.

Joyner:2008:OSC

[Joy08] David Joyner. Open source computer algebra systems: Axiom. ACM Communications in Computer Algebra, 42(1–2):39–47, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Jenks:1992:ASC

[JS92] Richard D. Jenks and Robert S. Sutor. AXIOM: The Scientific Computation System. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97855-0 (New York), 3-540-97855-0 (Berlin). xxiv + 742 pp. LCCN QA76.95.J46 1992.

URL http://link.springer.com/
10.1007/978-1-4612-2940-7.

Jenks:2003:AVS

[JSD03] Richard D. Jenks, Robert S. Sutor, and Tim Daly. Axiom Volume 0: The Scientific Computation System. 2003. xviii + 1187 pp. LCCN

????  URL http://www.axiom-
developer.org/axiom-website/
bookvol0.pdf.

Jenks:1986:SIA

[JSW86] Richard D. Jenks, Robert S. Sutor, and Stephen M. Watt. Scratchpad II: an abstract datatype system for mathematical computation. Research Report RC 12327 (#55257), International Business Machines Inc., Thomas J. Watson Research Center, Yorktown Heights, NY, USA, 1986. 23 pp.

Jenks:1988:SIA

[JSW88] R. D. Jenks, R. S. Sutor, and S. M. Watt. Scratchpad II: An abstract datatype system for mathematical computation. In Janßen [Jan88], pages 12–37. ISBN 3-540-18928-9, 0-387-18928-9. LCCN QA155.7.E4T74 1988.

Jenks:1994:HMA

[JT94] Richard D. Jenks and Barry M. Trager. How to make AXIOM into a Scratchpad. In ACM [ACM94], pages 32–40. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994.

Kendall:2001:SIC

[Ken01] Wilfrid S. Kendall. Symbolic Itô calculus in AXIOM: an ongoing story. Statistics and Computing, 11(1):25– 35, 2001. CODEN STACE3. ISSN 0960-3174.

Kusche:1989:IGT

[KKM89] K. Kusche, B. Kutzler, and H. Mayr. Implementation of a geometry theorem proving package in SCRATCHPAD II. In Davenport [Dav89], pages 246– 257. ISBN 3-540-51517-8. LCCN QA155.7.E4E86 1987.

Keady:1994:PAS

[KN94] G. Keady and G. Nolan. Production of Argument SubPrograms in the AXIOM — NAG link: examples involving nonlinear systems. Technical Report TR1/94 ATR/7 (NP2680), Numerical Algorithms Group, Inc., Downer’s Grove, IL, USA and Oxford, UK, 1994. ?? pp.

URL http://www.nag.co.uk/doc/
TechRep/axiomtr.html.

Koseleff:1991:WGF

[Kos91] P.-V. Koseleff. Word games in free Lie algebras: several bases and formulas. Theoretical Computer Science, 79(1):241–256, February 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lambe:1991:RHP

[Lam91] L. A. Lambe. Resolutions via homological perturbation. Journal of Symbolic Computation, 12(1):71–87, July 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Lambe:1994:NGC

[Lam94] Larry Lambe. Next generation computer algebra systems AX-IOM and the Scratchpad concept: Applications to research in algebra. In Mats Gyllenberg and Lars Erik Persson, editors, Analysis, algebra, and computers in mathematical research: proceedings of the Twenty-first Nordic Congress of Mathematicians, Luleå University of Technology, Sweden, 1992, volume 156 of Lecture Notes in Pure and Applied Mathematics, pages 201–222. Marcel Dekker, New York, NY, USA, 1994. ISBN 0-8247-9217-3. LCCN QA299.6 .N67 1992. URL http://www.

loc.gov/catdir/enhancements/
fy0647/94002464-d.html.

LeBlanc:1991:UMT

[LeB91] S. E. LeBlanc. The use of MathCAD and Theorist in the ChE classroom. In Anonymous [Ano91], pages 287– 299 (vol. 1). 2 vol.

Levelt:1995:IPI

[Lev95] A. H. M. Levelt, editor. ISSAC ’95: Proceedings of the 1995 International Symposium on Symbolic and Algebraic Computation: July 10–12, 1995, Montréal, Canada, IS-SAC -PROCEEDINGS- 1995. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-699-9. LCCN QA 76.95 I59 1995. ACM order number: 505950.

Lynch:1991:NQM

[LM91] R. Lynch and H. A. Mavromatis. New quantum mechanical perturbation technique using an ‘electronic scratchpad’ on an inexpensive computer. American Journal of Physics, 59(3):270–273, March 1991. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Li:2006:EIP

[LM06] Xin Li and Moreno Maza. Efficient implementation of polynomial arithmetic in a multiple-level programming environment. Lecture Notes in Computer Science, 4151: 12–23, 2006. CODEN LNCSD9. ISBN 3-540-38084-1. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://www.csd.
uwo.ca/~moreno//Publications/
Li-MorenoMaza-ICMS-06.pdf. Pro-

ceedings of the International Congress of Mathematical Software (ICMS 2006).

Li:2007:VGP

[LMS07] Xin Li, Marc Moreno Maza, and Éric Schost. On the virtues of generic programming for symbolic computation. In Shi et al. [SvADS07], pages 251–258. CO-DEN LNCSD9. ISBN 3-540-72585-7 (print), 3-540-72586-5 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.
com/content/978-3-540-72586-2.

Lucks:1986:FIP

[Luc86] Michael Lucks. A fast implementation of polynomial factorization. In Bruce W. Char, editor, Proceedings of the 1986 Symposium on Symbolic and Algebraic Computation: Symsac ’86, July 21–23, 1986, Waterloo, Ontario, pages 228–232. ACM Press, New York, NY 10036, USA, 1986. ISBN 0-89791-199-7. LCCN QA155.7.E4 A281 1986. ACM order number 505860.

Lueken:1977:UIF

[Lue77] E. Lueken. Ueberlegungen zur Implementierung eines Formelmanipulationssystemes. Master’s thesis, Technischen Universität Carolo-Wilhelmina zu Braunschweig, Braunschweig, Germany, 1977. ?? pp.

Mathews:1989:SCA

[Mat89] J. Mathews. Symbolic computational algebra applied to Picard iteration. Mathematics and computer education, 23(2):117–122, Spring 1989. CODEN MCEDDA. ISSN 0730-8639.

Miola:1990:DIS

[Mio90] A. Miola, editor. Design and Implementation of Symbolic Computation Systems, International Symposium DISCO ’90, Capri, Italy, April 10– 12, 1990, Proceedings, volume 429 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 0-387-52531-9 (New York), 3-540-52531-9 (Berlin). LCCN QA76.9.S88I576 1990.

Miola:1993:DIS

[Mio93] A. Miola, editor. Design and Implementation of Symbolic Computation Systems International Symposium. DISCO ’93 Gmunden, Austria, September 15–17, 1993: Proceedings. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-57235-X. LCCN QA76.9.S88I576 1993.

Monagan:1993:GPD

[Mon93] M. B. Monagan. Gauss: a parameterized domain of computation system with support for signature functions. In Miola [Mio93], pages 81– 94. ISBN 3-540-57235-X. LCCN QA76.9.S88I576 1993.

Mora:1989:AAA

[Mor89] T. Mora, editor. Applied Algebra, Algebraic Algorithms and Error-Correcting Codes. 6th International Conference, AAECC-6, Rome, Italy, July 4–8, 1988. Proceedings, volume 357 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-51083-4. LCCN QA268 .A35 1988. Conference held jointly with ISSAC ’88.

Melachrinoudis:1990:TAT

[MR90] E. Melachrinoudis and D. L. Rumpf.

Teaching advantages of transparent computer software — MathCAD. CoED, 10(1):71–76, January–March 1990. CODEN CWLJDP. ISSN 0736-8607.

Norman:1975:CFP

[Nor75] A. C. Norman. Computing with formal power series. ACM Transactions on Mathematical Software, 1 (4):346–356, December 1975. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Ollivier:1989:IRM

[Oll89] F. Ollivier. Inversibility of rational mappings and structural identifiability in automatics. In ACM [ACM89], pages 43–54. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

Online:1972:OCP

[Onl72] Online 72: conference proceedings … international conference on online interactive computing, Brunel University, Uxbridge, England, 4–7 September 1972. Online Computer Systems Ltd, Uxbridge, England, September 1972. ISBN 0-903796-02-3. LCCN QA76.55 .O54 1972. Two volumes.

Page:2007:AOS

[Pag07] William S. Page. Axiom: open source computer algebra system. ACM Communications in Computer Algebra, 41(3):114, September 2007. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Petrick:1971:PSS

[Pet71] S. R. Petrick, editor. Proceedings of the second symposium on Symbolic and Algebraic Manipulation, March 23–25, 1971, Los Angeles, California. ACM Press, New York, NY 10036, USA, 1971. LCCN QA76.5 .S94 1971.

Petitot:1993:EA

[Pet93] M. Petitot. Experience with Axiom. In Jacob et al. [JOS93], page 240.

Portes:2007:AVA

[PRDD07] Alfredo Portes, Arthur Ralfs, Timothy Daly, and Martin Dunstan. Axiom Volume 11: Axiom Browser. 2007. xix + 1193 pp. LCCN

????  URL http://www.axiom-
developer.org/axiom-website/
bookvol11.pdf.

Purtilo:1986:ASI

[Pur86] J. Purtilo. Applications of a software interconnection system in mathematical problem solving environments. In Bruce W. Char, editor, Proceedings of the 1986 Symposium on Symbolic and Algebraic Computation: Symsac ’86, July 21– 23, 1986, Waterloo, Ontario, pages 16–23. ACM Press, New York, NY 10036, USA, 1986. ISBN 0-89791-199-7. LCCN QA155.7.E4 A281 1986. ACM order number 505860.

Rioboo:1992:RAC

[Rio92] Renaud Rioboo. Real algebraic closure of an ordered field, implementation in Axiom. In Wang [Wan92], pages 206–215. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992.

Roesner:1995:VSP

[Roe95] K. G. Roesner. Verified solutions for parameters of an exact solution for non-Newtonian liquids using computer algebra. Zeitschrift fur Angewandte Mathematik und Physik, 75 (suppl. 2):S435–S438, 1995. ISSN 0044-2267.

Salvy:1989:EAA

[Sal89] B. Salvy. Examples of automatic asymptotic expansions. Technical Report 114, Inst. Nat. Recherche Inf. Autom., Le Chesnay, France, December 1989. 18 pp.

Salvy:1991:EAA

[Sal91] B. Salvy. Examples of automatic asymptotic expansions. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(2):4–17, April 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Schwarz:1988:PAD

[Sch88] F. Schwarz. Programming with abstract data types: the symmetry package SPDE in Scratchpad. In Janßen [Jan88], pages 167–176. ISBN 3-540-18928-9, 0-387-18928-9. LCCN QA155.7.E4T74 1988.

Schwarz:1989:FAL

[Sch89] F. Schwarz. A factorization algorithm for linear ordinary differential equations. In ACM [ACM89], pages 17–25. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989.

Schwarz:1991:MOG

[Sch91] F. Schwarz. Monomial orderings and Gröbner bases. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(1):10–23, January 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Smith:2007:ADA

[SDJ07] Jacob Smith, Gabriel Dos Reis, and Jaakko Järvi. Algorithmic differentiation in Axiom. In Brown [Bro07], pages 347–354. ISBN 1-59593-743-9 (print), 1-59593-742-0 (CD-ROM). LCCN QA76.5 S98 2007. ACM order number 505070.

Seiler:1994:CIA

[Sei94a] W. M. Seiler. Completion to involution in AXIOM. In Calmet [Cal94], pages 103–104.

Seiler:1994:PDO

[Sei94b] Werner M. Seiler. Pseudo differential operators and integrable systems in AXIOM. Computer Physics Communications, 79(2): 329–340, April 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.com/
science/article/pii/0010465594900760.

Sit:1989:GAS

[Sit89] W. Y. Sit. On Goldman’s algorithm for solving first-order multinomial autonomous systems. In Mora [Mor89], pages 386–395. ISBN 3-540-51083-4. LCCN QA268 .A35 1988. Conference held jointly with ISSAC ’88.

Sit:1992:ASP

[Sit92] W. Y. Sit. An algorithm for solving parametric linear systems. Journal of Symbolic Computation, 13 (4):353–394, April 1992. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Sutor:1987:TICb

[SJ87a] R. S. Sutor and R. D. Jenks. The type inference and coercion facilities in the Scratchpad II interpreter. In Wexelblat [Wex87], pages 56–63. ISBN 0-89791-235-7. LCCN QA76.7 .S54 v.22:7. US$23.00. SIGPLAN Notices, vol. 22, no. 7 (July 1987).

Sutor:1987:TICa

[SJ87b] Robert S. Sutor and Richard D. Jenks. The type inference and coercion facilities in the Scratchpad II interpreter. Research Report RC 12595 (#56575), IBM Thomas J. Watson Research Center, Yorktown Heights, NY, USA, 1987. 11 pp.

Smedley:1992:UPO

[Sme92] Trevor J. Smedley. Using pictorial and object oriented programming for computer algebra. In Hal Berghel et al., editors, Applied computing — technological challenges of the 1990’s: proceedings of the 1992 ACM/SIGAPP Symposium on Applied Computing, Kansas City Convention Center, March 1–3, 1992, pages 1243–1247. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-502-X. LCCN QA76.76.A65 S95 1992.

Shannon:1988:UGB

[SS88] D. Shannon and M. Sweedler. Using Gröbner bases to determine algebra membership, split surjective algebra homomorphisms determine birational equivalence. Journal of Symbolic Computation, 6(2-3):267– 273, October–December 1988. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Senechaud:1987:SIP

[SSV87] P. Senechaud, F. Siebert, and G. Villard. Scratchpad II: Présentation d’un nouveau langage de calcul formel. Technical Report 640-M, TIM 3 (IMAG), Grenoble, France, February 1987. ?? pp.

Sutor:1985:SIC

[Sut85] R. S. Sutor. The Scratchpad II computer algebra language and system. In Buchberger and Caviness [BC85], pages 32–33. ISBN 0-387-15983-5 (vol. 1), 0-387-15984-3 (vol. 2). LCCN QA155.7.E4 E86 1985. Two volumes.

Shi:2007:CSIb

[SvADS07] Yong Shi, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational Science — ICCS 2007: 7th International Conference, Beijing, China, May 27 — 30, 2007, Proceedings, Part II, volume 4488 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNCSD9. ISBN 3-540-72585-7 (print), 3-540-72586-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.
com/content/978-3-540-72586-2.

Sutor:1988:SIA

[SW88] R. D. Jenks R. S. Sutor and S. M.

Watt. Scratchpad II: An abstract datatype system for mathematical computation. In Janßen [Jan88], pages 12–?? ISBN 3-540-18928-9, 0-387-18928-9. LCCN QA155.7.E4T74 1988.

vanHoeij:1994:ACI

[vH94] M. van Hoeij. An algorithm for computing an integral basis in an algebraic function field. Journal of Symbolic Computation, 18(4):353–363, October 1994. CODEN JSYCEH.

ISSN 0747-7171 (print), 1095-855X (electronic).

Wang:1989:PCL

[Wan89] D. Wang. A program for computing the Liapunov functions and Liapunov constants in Scratchpad II. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 23(4):25–31, October 1989. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Wang:1991:MMC

[Wan91] Dongming Wang. Mechanical manipulation for a class of differential systems. Journal of Symbolic Computation, 12(2):233–254, August 1991. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Wang:1992:ISS

[Wan92] Paul S. Wang, editor. International System Symposium on Symbolic and Algebraic Computation 92. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-489-9 (soft cover), 0-89791-490-2 (hard cover). LCCN QA76.95.I59 1992.

Watt:1989:FPM

[Wat89] S. M. Watt. A fixed point method for power series computation. In Gianni [Gia89], pages 206–217. ISBN 3-540-51084-2. LCCN QA76.95 .I57 1988. Conference held jointly with AAECC-6.

Watt:1991:PIS

[Wat91] Stephen M. Watt, editor. Proceedings of the 1991 International Symposium on Symbolic and Algebraic Computation, ISSAC’91, July 15– 17, 1991, Bonn, Germany. ACM Press, New York, NY 10036, USA, 1991. ISBN 0-89791-437-6. LCCN QA76.95.I59 1991.

Weber:1993:CCA

[Web93] A. Weber. On coherence in computer algebra. In Miola [Mio93], pages 95–106. ISBN 3-540-57235-X. LCCN QA76.9.S88I576 1993.

Wexelblat:1987:IIT

[Wex87] Richard L. Wexelblat, editor. Proceedings of the SIGPLAN ’87 Symposium on Interpreters and Interpretive Techniques, St. Paul, Minnesota, June 24–26, 1987. ACM Press, New York, NY 10036, USA, 1987. ISBN 0-89791-235-7. LCCN QA76.7 .S54 v.22:7. US$23.00. SIG-PLAN Notices, vol. 22, no. 7 (July 1987).

Watanabe:1990:IPI

[WN90] Shunro Watanabe and Morio Nagata, editors. ISSAC ’90. Proceedings of the International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 1990. ISBN 0-89791-401-5. LCCN QA76.95 .I57 1990.

Zenger:1992:GFD

[Zen92] Ch. Zenger. Gröbnerbasen für Differentialformen und ihre Implementierung in AXIOM. Diplomarbeit, Universität Karlsruhe, Karlsruhe, Germany, 1992. ?? pp.